OpenStudy (anonymous):

The length of a football field is 67 yards greater than the width. The perimeter of the field is 346 yards. Find the length and width of the football field.

OpenStudy (anonymous):

L=212/4 L=53 YDS. IS THE LENGTH. W=53+67=120 YDS. FOR THE WIDTH. PROOF: 2*53+2*120=346 106+240=346 346=346

OpenStudy (anonymous):

Thnx sooo much sammie

OpenStudy (anonymous):

This is a bit of an algebra problem. First it's easiest to draw a picture:  We know the perimetre is 346, this means that \[2a+2b=346\]but we are also told that the length is 67 yards longer than the width, so we get \[b=a+67\] Now we can substitute this into the first equation we got. \[2a+2(a+67)=346\]\[2a + 2a + 134=346\]\[4a=212\]\[a=53\] Now we have a, we can go back and get b from our expression. \[b=53+67\]\[b=120\]So our final answer is that a=53, b=120.